OpenStudy (anonymous):

This is my last question before i finish my semester. Please help! Part 1: Using complete sentences, explain how to find the equation of the line, in standard form and slope–intercept form, passing through (–1, 5) and (1, 9). Part 2: Compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope–intercept form.

jimthompson5910 (jim_thompson5910):

first you need the slope of the line through (–1, 5) and (1, 9)

jimthompson5910 (jim_thompson5910):

use the formula m = (y2 - y1)/(x2 - x1)

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now use y = mx+b

jimthompson5910 (jim_thompson5910):

y = mx+b y = 2x + b ... plug in the slope 5 = 2(-1) + b ... plug in the given point (-1, 5) 5 = -2 + b ... ... ... b = ??

OpenStudy (anonymous):

7 ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so the equation in slope intercept form is y = 2x + 7

jimthompson5910 (jim_thompson5910):

to convert to standard form, we just get it into Ax+By = C form

jimthompson5910 (jim_thompson5910):

some books/teachers require A to be positive, so it's best to make A positive

jimthompson5910 (jim_thompson5910):

y = 2x + 7 y - 2x = 2x + 7 - 2x -2x + y = 7 2x - y = -7 The equation is now in standard form Ax+By = C where A > 0

jimthompson5910 (jim_thompson5910):

I'll leave part 2 to you

OpenStudy (anonymous):

leave what to me?

OpenStudy (anonymous):

oh okay i read it wrong

jimthompson5910 (jim_thompson5910):

that's ok